If one photon of light is known to have energy 3.33×10^-19J, is it visible?


1
Expert's answer
2021-05-16T17:59:26-0400

From the definition of the electron volt, we know 1 eV = 1.60 × 10−19 J

"E = \frac{3.33 \times 10^{-19}}{1.60 \times 10^{-19}} = 2.08 \;eV \\

λ = \frac{hc}{E} \\

= \frac{1.24 \times 10^{-6}}{2.08} = 596 \;nm"

Visible light: λ=400−700 nm

Answer: It is visible.